Circular No. 7358

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000G IN UGC 1773 T. Puckett, Mountain Town, GA, reports his discovery of an apparent

supernova (mag 17.2) on an unfiltered CCD frame (limiting mag 20.1) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Feb. 5.04 UT. SN 2000G is located at α = 2^{h}18^{m}15^{s}.27, δ = +13^{o}12′19′′.4 (equinox 2000.0), which is 0′′.13 east and 3′′.3 north of the center of UGC 1773. The new object was also present on an unfiltered CCD frame taken on Feb. 7.00, but it was not present on a frame taken on 1999 Sept. 7.34 — nor does SN 2000G appear on a Palomar Sky Survey images taken on 1990 Oct. 14 (limiting mag ∼ 21.2) or 1950 Aug. 14 (limiting mag ∼ 20.0).

P. Garnavich, University of Notre Dame; and S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000G was obtained by K. Rines on Feb. 7.2 UT with the F. L. Whipple Observatory 1.5-m telescope, which shows that it is a type-II supernova just past maximum light. A broad Hα emission feature is present, as well as Hβ absorption, which indicates a photospheric velocity of 8300 km/s. Narrow emission features from the host galaxy provide a recession velocity of 3600 km/s.

SAX J1808.4–3658 M. van der Klis, University of Amsterdam; D. Chakrabarty, J. C. Lee,

E. H. Morgan, and R. Wijnands, Massachusetts Institute of Technology; C. B. Markwardt, University of Maryland and Goddard Space Flight Center (GSFC); and J. H. Swank, GSFC, write: “The x-ray transient SAX J1808.4–3658, the only known accretion-powered millisecond pulsar (IAUC 6876, 6877), has been detected at a low x-ray-flux level of 3–12 mCrab (2– 10 keV) since Jan. 21.1 UT with the PCA instrument on the Rossi X-Ray Timing Explorer (RXTE). The start time of this outburst is not known because the source was unobservable by RXTE from ∼ 1999 Nov. 2 to 2000 Jan. 21, due to its proximity to the sun. On 1999 Nov. 2, the upper limit on the source flux was < 0.5 mCrab (2–10 keV). X-ray pulsations at the 401-Hz spin frequency were detected in several short observations between Jan. 21.1 and Feb. 6.2. Since Feb. 2, the source has exhibited violent quasi-periodic x-ray flaring with an rms amplitude varying between 40 and 100 percent of the average flux and a repetition frequency varying between 0.9 and 1.5 Hz. This is unlike any phenomenon previously observed in a neutron-star low-mass x-ray binary. The source is still close to the sun, but observations at other wavelengths (particularly radio and infrared) are strongly encouraged.”

2000 February 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green